# Search for gravitationally lensed interstellar transmissions

### v3

@@ Abstract
We consider interstellar light transmission facilitated by gravitational lensing. We study axially-symmetric lensing configurations where the transmitter, lens, and receiver are nearly aligned. With the transmitter positioned in the focal region of the lens, we investigate the caustic formed by a diffraction-limited beam of light emitted by the transmitter. We evaluated the impact of the lens’s point spread function (PSF) on the received beam. We estimate the power delivered to a receiver at interstellar distances, evaluate the major noise sources, and determine the detection sensitivity in both the noise- and signal-dominated regimes. Considering realistic assumptions about the transmitter’s performance, we explore signal detection strategies enhanced by the spatial broadening of the received beam, a result of the PSF from the transmitting lens. We find that detection of the lensed optical signals from nearby stars may be done relying on established optical engineering technologies. Furthermore, a network of astronomical facilities capable of observations in multiple narrow spectral bands will benefit the search. Our results support the feasibility of interstellar power transmission via gravitational lensing, directly contributing to the ongoing optical SETI efforts.
@@ II EM waves in a gravitational field
We consider a nearby star with mass $M_{\tt L}$, radius $R_{\tt L}$, and its Schwarzschild radius of $r_{g}=2GM_{\tt L}/c^{2}$. We assume a transmitter placed on the optical axis—the line connecting the solar system and this stellar gravitational lens, as illustrated in Fig. 1. We also assume that this transmitter is facing the lens and is positioned in the lens’ focal region at a distance $z_{0}>R^{2}_{\tt L}/2r_{g}=547.8\,(R_{\tt L}/R_{\odot})^{2}(M_{\odot}/M_{\tt L})$ AU from it, on the side opposite the solar system. Next, we assume that this is a CW laser optical transmitter that is capable of coherently illuminating an annulus around the lens with a radius of $b=\sqrt{2r_{g}z_{0}}>R_{\tt L}$, which is chosen to ensure the beam’s partial focusing at the receiver Turyshev (2024). To achieve this, a diffraction-limited transmitter must produce an annular beam of light. [^2] See Duocastella and Arnold (2012); Livneh et al. (2018) on the current efforts in producing annular laser beams. This beam should have the mean diverging angle of $b/z_{0}\simeq 7.16\times 10^{-6}\,({b}/{R_{\tt L}})(R_{\tt L}/{R_{\odot}})({650\,{\rm AU}}/{z_{0}})~{}{\rm rad}$ relative to the optical axis. As shown by Fig. 1, there is a finite-size caustic that is formed in the focal region of the lens. What determines the dimensions of this caustic? In the following discussion, we examine the geometry of the region where beam focusing occurs.
@@ II.1 Lensing caustic
Clearly, the rays with different impact parameters will intersect the optical axis at different distances, which is consistent with lens’ spherical aberration. As a result, for an infinitely large incident plane wave coming from infinity, a monopole gravitational lens forms a caustic with a semi-infinite focal line that begins from $z_{0}=R_{\tt L}/2r_{g}$ Turyshev and Toth (2017).
@@ II.2 Light amplification
Amplification (×$\times 10^{10}$) Refer to caption Distance from the optical axis, $\rho$ [m]
@@
II.3 Power projection by an extended source
@@ IV Detection strategies
Detecting laser signals presents several challenges. The first hurdle is the highly monochromatic nature of laser light, which emits at a specific wavelength. This characteristic necessitates pinpointing the exact wavelength, $\lambda$, to match an extraterrestrial signal and selecting a suitable narrow band-pass filter width, $\Delta\lambda$, to enhance the SNR, particularly in noise-dominated environments (36). Another challenge is the directionality of laser transmissions. Unlike radio waves, which can spread in all directions, laser beams are highly focused and narrow. Thus, for interstellar laser signals to be detected, they must be precisely aimed towards Earth, especially since interstellar gas and dust, though largely transparent to near-IR light, do not alleviate the need for exact alignment. Consequently, while the stellar brightness (34) and optical background noise are relatively well understood for each nearby star, the specific link parameters of the transmitted signal (32)—such as transmit power, aperture, wavelength, and pointing—are not known in advance.
Gravitational lensing makes some challenges above to be less critical, but some will require consideration: Although the beam of light we are searching for is significantly brighter compared to that of free-space laser transmissions, the transmitted beam arrives in the solar system being highly compressed. Clearly, this feature must be accounted for and appropriately addressed at the transmitter, thus to enhance the probability of signal detection.
@@ IV.1 Considering a search campaign
$\tau_{\tt beam}=\frac{d_{\tt beam}}{v_{\oplus}}\simeq 0.11\,\Big{(}\frac{z}{10\,{\rm pc}}\Big{)}\Big{(}\frac{650\,{\rm AU}}{z_{0}}\Big{)}\Big{(}\frac{d_{\tt T}}{1~{}{\rm m}}\Big{)}\Big{(}\frac{30~{}{\rm km/s}}{v_{\oplus}}\Big{)}~{}~{}{\rm s}.$ (39)
@@ IV.2 Plausible detection strategies
We note that directing an interstellar signal at a specific point in Earth’s heliocentric orbit could lead to false negatives. In such instances, a persistent signal, correlating with Earth’s orbital period, might be observed annually. If this signal is sufficiently weak, it could be mistaken for background noise and disregarded. Therefore, it is critical to observe the targets across multiple spectral bands. Detecting periodic anomalous photometric behavior in one or more channels—especially if multiple laser wavelengths are employed but not uniformly—could signal the presence of the signal of interest. Furthermore, a network of spatially distributed telescopes could facilitate early alerts and enable independent measurements, especially as they will sample the data with different values of factor $\mu(\rho_{0})$.
Once the signal is detected, she spatial distribution of receivers is invaluable, as each will capture a distinct dataset by traveling through the signal along a different path, as encoded $\mu(\rho_{0})$. Correlating the photometric and spectral data from each path enables the reconstruction of the beam’s full profile as it projected onto the solar system. Integrating this information with spectral data from multiple channels reveals the transmitter’s specific features encoded in the beam, such as its power, shape, design, and propulsion capabilities. Additionally, if the optical signal contains encoded information, transmitted via a set of particular patterns, this information will become accessible as well.
@@ V Discussion
Every star is destined to form an Einstein ring. This ring may form naturally with distant sources or through advanced technology, deliberately deployed to send us a message. The position of an Einstein ring around a star, as detailed in (20), depends on our distance to the star, the star’s mass, and the transmitter’s location. Although we can not resolve the rings that may form around nearby stellar lenses by light transmitted to us from their focal regions, with this information, we can initiate our search for the laser transmissions enabled by gravitational lensing.
@@
In the early phase of this process, a network of astronomical ground- and space-based astronomical facilities may provide valuable spatial, spectral and temporal information as microlensing events unfold. This data may be processed with existing tools Turyshev and Toth (2023) to recover the properties of the signal and potentially the information encoded within.

### v4

@@ Abstract
We explore interstellar light transmission facilitated by gravitational lensing, focusing on axially-symmetric lensing configurations where the transmitter, lens, and receiver are nearly aligned. Positioning an optical transmitter in the lens’s focal region, we investigate the caustic formed by a diffraction-limited annular beam of light emitted by the transmitter. We analyze the impact of the lens’s point spread function (PSF) on the projected beam’s structure, estimate the power delivered to a receiver at interstellar distances, and assess the major noise sources. We determine detection sensitivity in both noise- and signal-dominated regimes. Considering realistic assumptions about the transmitter’s performance, we explore signal detection strategies enhanced by the spatial broadening of the received beam, a result of the transmitting lens’s PSF. Our findings indicate that detecting lensed optical signals from nearby stars is achievable using established optical engineering technologies. A network of spatially distributed astronomical facilities capable of observations in multiple narrow spectral bands will enhance the search. Our results support the feasibility of interstellar power transmission via gravitational lensing, directly contributing to ongoing optical SETI efforts.
@@ II EM waves in a gravitational field
We consider a nearby star with mass $M_{\tt L}$, radius $R_{\tt L}$, and its Schwarzschild radius of $r_{g}=2GM_{\tt L}/c^{2}$. We assume a transmitter placed on the optical axis—the line connecting the solar system and this stellar gravitational lens, as illustrated in Fig. 1. We also assume that this transmitter is facing the lens and is positioned in the lens’ focal region at a distance $z_{0}>R^{2}_{\tt L}/2r_{g}=547.8\,(R_{\tt L}/R_{\odot})^{2}(M_{\odot}/M_{\tt L})$ AU from it Lodge (1919); Chwolson (1924); Einstein (1936); Eshleman (1979); Turyshev and Andersson (2003); Turyshev and Toth (2017), on the side opposite the solar system. Next, we assume that this is a CW laser optical transmitter that is capable of coherently illuminating an annulus around the lens with a radius of $b=\sqrt{2r_{g}z_{0}}>R_{\tt L}$, which is chosen to ensure the beam’s partial focusing at the receiver Turyshev (2024). To achieve this, a diffraction-limited transmitter must produce an annular beam of light. [^2] See Duocastella and Arnold (2012); Livneh et al. (2018) on the current efforts in producing annular laser beams. This beam should have the mean diverging angle of $b/z_{0}\simeq 7.16\times 10^{-6}\,({b}/{R_{\tt L}})(R_{\tt L}/{R_{\odot}})({650\,{\rm AU}}/{z_{0}})~{}{\rm rad}$ relative to the optical axis. As shown by Fig. 1, there is a finite-size caustic that is formed in the focal region of the lens. What determines the dimensions of this caustic? In the following discussion, we examine the geometry of the region where beam focusing occurs.
@@ II.1 Lensing caustic
Clearly, the rays with different impact parameters will intersect the optical axis at different distances, which is consistent with lens’ spherical aberration. As a result, for an infinitely large incident plane wave coming from infinity, a monopole gravitational lens forms a caustic with a semi-infinite focal line that begins from $z_{0}=R_{\tt L}/2r_{g}$.
@@ II.2 Light amplification
Amplification (×$\times 10^{10}$)
Distance from the optical axis, $\rho$ [m]
@@
II.3 Power projected by a stellar gravitational lens
@@ IV Detection strategies
Detecting laser signals presents several challenges. The first hurdle is the highly monochromatic nature of laser light, which emits at a specific wavelength. This characteristic necessitates pinpointing the exact wavelength, $\lambda$, to match an extraterrestrial signal and selecting a suitable narrow band-pass filter width, $\Delta\lambda$, to enhance the SNR, particularly in noise-dominated environments (36). Another challenge is the directionality of laser transmissions. Unlike radio waves, which can spread in all directions, laser beams are highly focused and narrow. Thus, for interstellar laser signals to be detected, they must be precisely aimed towards the Earth, especially since the interstellar gas and dust, though largely transparent to near-IR light, do not alleviate the need for exact alignment. Consequently, while the stellar brightness (34) and optical background noise are relatively well understood for each star in our stellar neighborhood, the specific link parameters of the transmitted signal (32)—such as the transmitted power, aperture, wavelength, and pointing direction used for that purpose—are not known in advance.
Gravitational lensing makes some challenges above to be less critical, but some will require consideration: Although the beam of light we are searching for is significantly brighter compared to that of a free-space laser transmission, the transmitted beam arrives in the solar system being highly compressed. Clearly, this feature must be accounted for and be appropriately addressed at the transmitter, thus to enhance the probability of signal detection.
@@ IV.1 Considering a search campaign
$\tau_{\tt beam}=\frac{d_{\tt beam}}{v_{\oplus}}\simeq 0.11\,\Big{(}\frac{z}{10\,{\rm pc}}\Big{)}\Big{(}\frac{650\,{\rm AU}}{z_{0}}\Big{)}\Big{(}\frac{d_{\tt T}}{1~{}{\rm m}}\Big{)}\Big{(}\frac{30~{}{\rm km/s}}{v_{\oplus}}\Big{)}~{}{\rm s}.$ (39)
@@ IV.2 Plausible detection strategies
We note that directing an interstellar signal at a specific point in Earth’s heliocentric orbit could lead to false negatives. In such instances, a persistent signal, correlating with Earth’s orbital period, might be observed annually. If this signal is sufficiently weak, it could be mistaken for background noise and disregarded. Therefore, it is critical to observe the targets across multiple spectral bands. Detecting periodic anomalous photometric behavior in one or more channels—especially if multiple laser wavelengths are employed but not uniformly—could signal the presence of the signal of interest. Furthermore, a network of spatially distributed telescopes could facilitate early alerts and enable independent measurements, especially as they will sample the data with different values of the factor $\mu(\rho_{0})$.
Once the signal is detected, the spatial distribution of receivers is invaluable, as each will capture a distinct dataset by traveling through the signal along a different path, as given by $\mu(\rho_{0})$. Correlating the photometric and spectral data from each path enables the reconstruction of the beam’s full profile as it projected onto the solar system. Integrating this information with spectral data from multiple channels reveals the transmitter’s specific features encoded in the beam, such as its power, shape, design, and propulsion capabilities. Additionally, if the optical signal contains encoded information, transmitted via a set of particular patterns, this information will become accessible as well.
@@ V Discussion
Every star is destined to form an Einstein ring. This ring may form naturally with distant sources present in the background (see relevant discussion in Turyshev and Toth (2023)) or through advanced technology, deliberately deployed to send us a message, the case discussed in Turyshev (2024) and considered in this paper. The position of an Einstein ring around a star, as detailed in (20), depends on our distance to the star, the star’s mass, and the transmitter’s location. Although we can not resolve the rings that may form around nearby stellar lenses by light transmitted to us from their focal regions, with this information, we can initiate our search for the laser transmissions enabled by gravitational lensing.
@@
In the early phases of this process, a network of astronomical ground- and space-based astronomical facilities may provide valuable spatial, spectral and temporal information as microlensing events unfold. This data may be processed with existing tools Turyshev and Toth (2023) to recover the properties of the signal and potentially the information encoded within.